# A solution of the problem of standard compact Clifford-Klein forms

### v1

@@ 6.1 Tables
Non-maximal irreducible subalgebras
Type of inclusion dimension $\omega_{r}$
$A_{n}\subset A_{l}$ $3\binom{n+2}{4}$ 00
$A_{n}\subset A_{l}$ $3\binom{n+3}{4}$ 00
@@
Irreducible triples
Inclusion 1 Inclusion 2 $\dim$ of the representation
$B_{n}\subset A_{2n}$ $A_{2n}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n+1}{k}$
$D_{n}\subset A_{2n-1}$ $A_{2n-1}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n}{k}$
Table 4:
6.2 Explanations to Tables
In Table 2 $\bar{\omega}_{r}$ is 00, if the representation is linear, but neither symplectic, nor orthogonal, it is $1$ is the representation is orthogonal, and it is $(-1)$ if the representation is symplectic. Table 3 reproduces inclusions $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$ of complex simple Lie algebras when $\mathfrak{h}^{c}$ is not maximal in $\mathfrak{g}^{c}$ from [14], Table 1 (recall that we reproduce only the infinite series). Table 4 reproduces inclusions of triples of irreducible inclusions $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}\subset\tilde{\mathfrak{g}}^{c}$ together with the dimensions of $\tilde{\mathfrak{g}}^{c}$. Here we need to explain some simplification in Tables 4 and 5. In [14] the author is concerned with the classification of subalgebras up to conjugation. Then non-conjugate subalgebras differ by integer labels $k$ in the corresponding Dynkin diagrams. These numbers yield different values of $\dim\tilde{\mathfrak{g}}^{c}$. However, our argument does not use this more rained classification, since we consider all possible real ranks for all possible real forms. Therefore, we do not indicate different (non-conjugate) inclusions of the same subalgebra.
6.3 An estimate
Lemma 12.
The following inequality holds for $l\geq 3$
$\binom{2l}{r}-\binom{2l}{r-2}>l^{2},2\leq r\leq l.$ (4)
Proof.
For $r=2$ one obtains
$\binom{2l}{2}-1=\frac{(2l-1)2l}{2}-1=2l^{2}-l-1>l^{2},\,\text{for}\,l>1.$
For $r=3$ one obtains
$\binom{2l}{3}-\binom{2l}{1}=\frac{1}{3}l(4l^{2}-2l-\frac{4}{3}),\,\text{for}\,l\geq 3.$
For $r=4$ one obtains
$\binom{2l}{4}-\binom{2l}{2}=\frac{1}{12}l(2l-1)(4l^{2}-10l+6),\,\text{for}\,l\geq 4.$
For $r\geq 5$ we use the formula
$\binom{n}{k}=\prod_{i=1}^{k}\frac{n+1-i}{i}.$
In our case
$\binom{2l}{r}-\binom{2l}{r-2}=\prod_{i=1}^{r}\frac{2l+1-i}{i}-\prod_{i=1}^{r-2}\frac{2l+1-i}{i}=\binom{2l}{r-2}\cdot(\frac{(2l+2-r)(2l+1-r)}{(r-1)r}-1)=$
$=\binom{2l}{r-2}\cdot\frac{4l^{2}-4lr+6l-2r+2}{r(r-1)}\geq\binom{2l}{3}\cdot\frac{4l^{2}-4lr+6l-2r+2}{r(r-1)}=(*)$
Notice that (*)(*)( * ) obtains the minimal value for $r=l$ and so
(*$(*)\geq\binom{2l}{3}\cdot\frac{4l+2}{l(l-1)}>2l^{2}-3l+1>l^{2},\,\text{for}\,l>4.$

### v3

@@ 6 Tables
Maximal subalgebras of maximal rank in $\mathfrak{g}^{c}$
$\mathfrak{g}^{c}$ $\mathfrak{\tilde{h}}^{c}$ $\operatorname{rank}\mathfrak{\tilde{h}}^{c}$
$\mathfrak{so}(2l+1,\mathbb{C})$ $\mathfrak{so}(2k,\mathbb{C})\oplus\mathfrak{so}(2(l-k)+1,\mathbb{C})$ $l$
$\mathfrak{sp}(2l,\mathbb{C})$ $\mathfrak{sp}(2k,\mathbb{C}\oplus\mathfrak{sp}(2(l-k)+1,\mathbb{C})$ $l$
$\mathfrak{so}(2l,\mathbb{C})$ $\mathfrak{so}(2k,\mathbb{C})\oplus\mathfrak{so}(2(l-k)+1),\mathbb{C})$ $l$
$\mathfrak{sl}(n,\mathbb{C})$ $\mathfrak{so}(2,\mathbb{C})\oplus\mathfrak{so}(2l-1,\mathbb{C})$ $l-1$
$\mathfrak{so}(2l+2,\mathbb{C})$ $\mathfrak{so}(2,\mathbb{C})\oplus\mathfrak{so}(2l-1),\mathbb{C})$ $l-1$
$\mathfrak{sp}(2l,\mathbb{C})$ $\mathfrak{gl}(l,\mathbb{C})$ $l-1$
$\mathfrak{so}(2l,\mathbb{C})$ $\mathfrak{so}(2,\mathbb{C})\oplus\mathfrak{so}(2l-2,\mathbb{C})$ $l-1$
$\mathfrak{so}(2l,\mathbb{C})$ $\mathfrak{gl}(l,\mathbb{C})$ $l-1$